# AweDist: Attention-aware Embedding Distillation for New Input Token Embeddings

### v1

@@ 6 Conclusion
In this work, we have described a fundamental limitation of many existing embedding initialization methods, which only exploit knowledge stored in a model’s embedding matrices, whereas much of the knowledge about token compositions actually resides in the Transformer layers (Elhage et al., 2022; Lad et al., 2024). To address this limitation, we have proposed AweDist, which systematically incorporates this knowledge via distilling contextual information from hidden states. Experimental results confirm that our method not only clearly outperforms “simple” aggregation of embedding matrix rows but also yields better performance than training embeddings using traditional causal language modeling as well as hyper-networks extensively pretrained for new token embedding prediction (Minixhofer et al., 2024). In future work, we believe research into embedding initialization should move beyond an aggregation of information stored in the embedding tables. In particular, a more localized identification of subtoken contextualization (e.g., specific attention heads that aggregate subtokens) is a promising direction for a more targeted distillation objective.
@@ Acknowledgements
Appendix A Limitations
Unknown tokens. Since the motivation for the distillation-based objective is to match the original model’s behavior, our method is generally not applicable if the original model cannot meaningfully process a new token. Note that this specifically limits the applicability of our method for adapting a model’s vocabulary to a previously unseen language.
Only input embeddings. As discussed in Section 3, our method only addresses the task of inducing input embeddings, since our distillation objective is not applicable to output embeddings. Our method can be freely combined with any other method for obtaining valid output embeddings. Alternatively, new tokens can be added only in the input embeddings, since the Transformer architecture does not require the same input and output vocabulary (although this assumption is made in some popular implementations).
Short subword tokens. We find that our method does not work well for polysemic tokens, i.e., tokens with many different meanings in different contexts. These are often short subwords like <ed> or <fer> that occur in many words. This also means that our method is not ideal to initialize embeddings for an entirely new vocabulary that contains such new subwords. However, these tokens could be initialized using other methods [Minixhofer et al., 2024, Gee et al., 2022], while our method is applied to the rest.
@@ B.1 Token Selection for Domain Adaptation
Therefore, we select all whole words from the chosen benchmarks that are not already existing tokens in the original vocabulary. We exclude tokens that include digits or any of the characters in the following string in order to reduce noise: {minted}[ framesep=2mm, baselinestretch=1.2, fontsize=, linenos ]Python "[]()<>.,;:!?@#
B.2 Multi-Word Token Generation
We prompted gpt-o4-mini-high (as of April 7, 2025) with the following prompt: {minted}[fontsize=, breaklines=true, linenos]python f"""Provide a list of entities, places, people, concepts, phrases or phrasings, idioms, or other terms that span multiple words, at least two and maximum five. Provide them in JSON format grouped by category and at least 100 per category. Ensure that you do not shorten coherent names just to fit them into the five word limit.""" We refined this prompt to exclude common failure modes through trial-and-error. We include the list of resulting multi-word tokens in our GitHub repository.
B.3 Open Medical-LLM Leaderboard Experiments
@@
For evaluation, we use the lm-evaluation-harness [Biderman et al., 2024] library at the version 0.4.7. We use 5-shot evaluation on a single Nvidia H100 GPU using bfloat16 precision. Specifically, we use the following command:
[fontsize=, breaklines=true, linenos]sh lm_eval –model hf –model_args pretrained=
B.3.1 Compute Budget & Runtime
We report the average training time for AweDist as well as NTP++ for initializing the new biomedical domain tokens per model in Table 7. We also report the number of new tokens per model. These differ slightly, as we build an initial fixed candidate list, which is then filtered against the existing vocabularies of each model.
Additional computational budget.
Since AweDist includes an additional forward pass through the model for generating “teacher” hidden states, we do see this overhead reflected in the run times. A forward pass can be approximated as taking one third of the computational cost of a complete forward-backward pass – which roughly matches our measured run times. Note however that – as demonstrated in Section 5.2 – AweDist can be significantly sped up while potentially even improving results by choosing an earlier target layer than the last one, yielding even faster run times than NTP++.
However, our main version of AweDist is slightly more expensive than NTP++ – even though both have the same data budget. Thus, we also run the baseline NTP++ with an additional epoch to compensate for the additional teacher forward pass that AweDist uses. Note that this in fact significantly overcompensates for the additional budget AweDist uses. We report the results in Table 8. NTP++ is not able to take any major advantage of an additional epoch, yielding similar results to using a single epoch. We investigate the loss curves and find that the training converges towards the beginning of the second epoch.666Note that we do not anneal the learning rate to zero, so this is not an artifact of the learning rate schedule.
We believe that this is at least partly due to the fundamental limitation of next-token prediction for our task of learning a single new embedding for an already pretrained model. Instead of matching model behavior between seeing the single new and the multiple original subtokens, NTP simply adjusts weights to maximize the likelihood of the given sequences. Consider “<new_token> is a football player”, where we want <new_token> to represent Messi: NTP will only be able to learn a generic embedding from this sequence that will capture semantics from many different (American football and soccer) players. Our proposed distillation objective AweDist however is able to learn a more specific representation, as it has access to hidden states from the sequence “Messi is a football player”, which capture more specific details necessary for a good representation.
To demonstrate, we also run AweDist for an additional epoch. AweDist is in fact able to take advantage of a second pass over the data, in some case with major improvements in the benchmark results. This illustrates the richer training signal that our distillation-based objective provides. Remarkably, for OLMo-2-7B-i, AweDist now is actually able to match the original model with the original tokenization while instead using the new tokenization.
Init Method Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i Llama-3-8B Llama-3-8B-i Mistral-7B OLMo-2-7B-i NTP++ 08:44 08:49 05:52 05:18 08:43 08:46 06:01 07:40 AweDist 11:31 11:33 07:38 08:07 12:00 11:29 09:09 10:22 # new tokens 2589 2589 2589 2589 2589 2589 2637 2591
Mistral-7B Llama-3-8B Llama-3-8B-i Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i OLMo-2-7B-i Avg. Target 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 NTP++ (1 epoch) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 AweDist (1 epoch) 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 NTP++ (2 epoch) 61.0 66.1 65.4 65.9 68.7 56.7 61.5 58.9 63.0 AweDist (2 epoch) 63.2 68.5 68.9 67.7 71.8 58.4 63.6 61.3 65.4
AweDist versus hyper-networks.
Comparing the computational budget with ZeTT [Minixhofer et al., 2024], ZeTT sees 3.2 billion tokens during hyper-network pretraining, while AweDist sees only 3.2 million tokens (for initializing 2,600 new tokens), less than 1% of ZeTT’s budget – still AweDist outperforms ZeTT. In return, ZeTT is faster at inference time (in our experiments on a single H100 80GB GPU, ZeTT took less than a minute). In future work, we believe that investigating an AweDist-style objective for hyper-network pretraining would be beneficial, as our experiments show that this significantly outperforms next-token prediction for learning new embeddings (ZeTT uses next-token prediction).
B.4 LLM-as-a-Judge Experiments
@@
For generating definitions for a particular new token, we use the prompt “The word <new_token> is defined as” following Teehan et al. [2024], where <new_token> is replaced by the string representation of the new token (with whitespace stripped from the left side). We use greedy decoding. Our evaluation considers the model checkpoint corresponding to the best learning rate from the experiments described in Section B.3 for each method.
@@
LLM-as-a-Judge evaluations are not perfect – see for example in Table 2 the similarity scores (Sim) of the target model (measuring similarity with the target model), which should obviously always be 100%, but turn out to be 99.4% / 99.8% in the case of OLMo-2-7B-i and Mistral-7B, respectively. Nevertheless, these results (which we report as a sanity check on the judge quality) are very close to correct, which is encouraging.
@@
We use meta-llama/Llama-3.3-70B-Instruct as the judge model and run it using bfloat16 layer-parallel inference over two H100 GPUs using HuggingFace transformers. We use prompts adapted from Li et al. [2023] and Villegas et al. [2025] for evaluating the general correctness of generated definitions as well as the similarity with the original model.
@@
[fontsize=, breaklines=true, linenos]python f"""<|start_header_id|>system<|end_header_id|>
You are a pattern-following assistant that can only answer with "Yes" or "No". Your goal is to determine whether a provided definition for a given word is correct. The definition should be on topic and specific but does not need to be exhaustive.<|eot_id|><|start_header_id|>user<|end_header_id|>
Remember to answer with one word either "Yes" or "No".
### Instruction: Determine if the provided definition for the word "token.lstrip()" is correct.
### Definition token.lstrip(): line["model_definition"].strip()
[…]
@@
[fontsize=, breaklines=true, linenos]python f"""<|start_header_id|>system<|end_header_id|>
You are a pattern-following assistant that can only answer with "Yes" or "No". Your goal is to determine whether a predicted definition conveys a similar enough meaning to the ground truth definition provided for a given word.<|eot_id|><|start_header_id|>user<|end_header_id|>
Remember to answer with one word either "Yes" or "No".
### Instruction: Determine if the predicted definition conveys a similar meaning to the ground truth definition. The word is "token.lstrip()".
### Ground truth definition: line["target_definition"].strip()
### Predicted definition: line["model_definition"].strip()
[…]
B.5 Generation of Relevant Contexts
When generating contexts that contain a new token, we prompt the model with the beginning-of-sequence (BOS) token followed by the textual representation of the new token. Using the new token “palatable” as an example, our prompt is then: “<s> palatable”. Note that we add a prefix space after the BOS token <s>. We then sample new sequences of length , where and are the same as in the settings we use for context retrieval from an existing corpus (in this work we use and ). We sample using the standard sampling settings of each model from the transformers [Wolf et al., 2020] library.
B.6 Isolated Per-Token Optimization vs. Joint Optimization
When implementing AweDist, we have the choice of performing the optimization process for each new input embedding separately or simply allowing updates to all new input embeddings at the same time.
@@ B.7 Hyperparameters
We use AdamW [Kingma and Ba, 2017, Loshchilov and Hutter, 2019] optimization for all trainable parameters (the new token embeddings) with a batch size of 16, which maximizes throughput on our used Nvidia H100 80GB GPUs. We do not incorporate any weight decay and maintain a constant learning rate schedule with a linear warmup for the first half of the training steps. For fair comparison, we sweep for the best learning rate for each method that requires a learning rate. We use the same data for all training-based methods, including baselines and variations of our method.
B.8 Models
Appendix C Additional Results
We repeat the analysis of the best target layer for AweDist from Section 5.2 in Figure 3 for Llama-3.1-8B-Instruct instead of Llama-3-8B-Instruct. Our analysis for Llama-3-8B-Instruct in Section 5.2 applies.

### v2

@@ 6 Conclusion
In this work, we have described a fundamental limitation of many existing embedding initialization methods, which only exploit knowledge stored in a model’s embedding matrices, whereas much of the knowledge about token compositions actually resides in the Transformer layers (Elhage et al., 2022; Lad et al., 2024). To address this limitation, we have proposed Token Distillation, which systematically incorporates this knowledge via distilling contextual information from hidden states. Experimental results confirm that our method not only clearly outperforms “simple” aggregation of embedding matrix rows but also yields better performance than training embeddings using traditional causal language modeling as well as hyper-networks extensively pretrained for new token embedding prediction (Minixhofer et al., 2024). In future work, we believe research into embedding initialization should move beyond an aggregation of information stored in the embedding tables. In particular, a more localized identification of subtoken contextualization (e.g., specific attention heads that aggregate subtokens) is a promising direction for a more targeted distillation objective.
Limitations
Unknown tokens. Since the motivation for the distillation-based objective is to match the original model’s behavior, our method is generally not applicable if the original model cannot meaningfully process a new token. Note that this specifically limits the applicability of our method for adapting a model’s vocabulary to a completely unseen language, although in current pretraining regimes, many languages do end up being seen to a certain extent due to imperfect filtering mechanisms or deliberate inclusion.
Only input embeddings. As discussed in Section 3, our method only addresses the task of inducing input embeddings, since our distillation objective is not applicable to output embeddings. Our method can be freely combined with any other method for obtaining valid output embeddings. Alternatively, new tokens can be added only in the input embeddings, since the Transformer architecture does allow for input tokens which do not occur in the output vocabulary (although this assumption is made in some popular implementations).
Short subword tokens. We find that our method does not work well for lexically ambiguous tokens, i.e., tokens with many different meanings in different contexts. These are often short subwords like <ed> or <fer> that occur in many words. This also means that our method is not ideal to initialize embeddings for an entirely new vocabulary that contains such new subwords. However, these tokens could be initialized using other methods (e.g.Minixhofer et al., 2024), while our method is applied to the rest.
@@ Acknowledgements
Appendix A Additional Results
A.1 Results for “Tokens to Words” (Kaplan et al., 2025)
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP (tune all embeddings) 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 Tokens to Words (Kaplan et al., 2025) (61.7) (64.3) (65.0) (63.9) (66.6) (57.3) (61.5) (59.5) (62.5) NTP (Lampinen et al., 2018) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
Mistral-7B Llama3-8B Llama3-8B-i Qwen3-8B Avg. Original tokenization 69.5 69.4 72.1 81.7 73.2 Random (Hewitt, 2021) 51.0 46.1 51.5 62.1 52.7 Mean (Gee et al., 2022) 56.3 58.4 61.7 69.6 61.5 Tokens to Words (Kaplan et al., 2025) (64.8) (62.6) (66.6) (82.2) (69.1) NTP (Lampinen et al., 2018) 64.7 67.0 70.1 81.2 70.8 ZeTT (Minixhofer et al., 2024) 66.1 61.3† 65.2† – – Token Distillation 68.5 68.9 72.9 81.5 72.9
We report results for the method from “Tokens to Words” (Kaplan et al., 2025) in Table 8 and Table 9 alongside our method and other baselines. Note that the “Tokens to Words” method only initializes embeddings for tokens where a PatchScopes (Ghandeharioun et al., 2024) extraction using a special prompt to identify hidden layers is successful. It is difficult to apply their method if the extraction is not successful. To give the fairest assessment of the method, we do not add these “unsuccessful” tokens to the vocabulary when using “Tokens to Words”. However, this likely overestimates the method’s performance, since the model now uses more of the original embeddings during testing, which have been tuned during massive pretraining. On average on our biomedical domain adaptation results, roughly 10–20% of the total tokens did not have a successful PatchScopes extraction, varying across models (up to 35% in the case of Llama3.2-3B and 40% in the case of OLMo2-7B-i). On our French language adaptation experiments, this is even more severe, ranging from 50% to 75% of unsuccessful extractions.
A.2 Different Target Layers for Token Distillation
We repeat the analysis of the best target layer for Token Distillation from Section 5.2 in Figure 3 for Llama3.1-8B-Instruct instead of Llama3-8B-Instruct. Our analysis for Llama3-8B-Instruct in Section 5.2 applies.
@@ B.1 Token Selection for Domain Adaptation
Therefore, we select all whole words from the chosen benchmarks that are not already existing tokens in the original vocabulary. We exclude tokens that include digits or any of the characters in the following string in order to reduce noise:
Additionally, to exclude very rare words that would not be commonly used as new tokens, we apply the following two filters: (1) We exclude words that appear fewer than five times across all benchmarks, and (2) we exclude words that occur fewer than 25 times in a sample of 768,000 documents from a domain-specific reference corpus (we use ncbi/pubmed for biomedical benchmarks). The full list of new words is available in our GitHub repository. For illustrative purposes, we provide a random sample of 100 added tokens here:
B.2 Token Selection for French Language Adaptation
We use a similar approach as described in Section B.1 but use HuggingFaceFW/fineweb-2 as a domain-specific reference corpus for filtering rare words and add words that appear in our chosen benchmarks from FrenchBench (Faysse et al., 2025). We also provide a random sample of 100 added tokens for illustrative purposes here:
B.3 Multi-Word Token Generation
We prompted gpt-o4-mini-high (as of April 7, 2025) with the following prompt:
We refined this prompt to exclude common failure modes through trial-and-error. We include the list of resulting multi-word tokens in our GitHub repository.
B.4 Open Medical-LLM Leaderboard Experiments
@@
For evaluation, we use the lm-evaluation-harness (Biderman et al., 2024) library at the version 0.4.7. We use 5-shot evaluation on a single Nvidia H100 GPU using bfloat16 precision. Specifically, we use the following command:
This evaluates the task of the Open Medical-LLM Leaderboard (Pal et al., 2024; Jin et al., 2019; 2020; Hendrycks et al., 2021; Pal et al., 2022). All code and a lock-file with specific versions will be available in our GitHub repository.
B.5 FrenchBench Experiments
As for the Open Medical-LLM Leaderboard Experiments, we use the lm-evaluation-harness (Biderman et al., 2024) library at the version 0.4.7 and use 5-shot evaluation on a single Nvidia H100 GPU with bfloat16 precision. We evaluate on the multiple-choice tasks from FrenchBench (Faysse et al., 2025) and use letters (e.g., A, B, C or D) as the answer choices. Specifically, we use the following command:
B.6 LLM-as-a-Judge Experiments
@@
For generating definitions for a particular new token, we use the prompt “The word <new_token> is defined as” following Teehan et al. (2024), where <new_token> is replaced by the string representation of the new token (with whitespace stripped from the left side). We use greedy decoding. Our evaluation considers the model checkpoint corresponding to the best learning rate from the experiments described in Section B.4 for each method.
@@
LLM-as-a-Judge evaluations are not perfect – see for example in Table 2 the similarity scores (Sim) of the target model (measuring similarity with the target model), which should obviously always be 100%, but turn out to be 99.4% / 99.8% in the case of OLMo2-7B-i and Mistral-7B, respectively. Nevertheless, these results (which we report as a sanity check on the judge quality) are very close to correct, which is encouraging.
@@
We use meta-llama/Llama3.3-70B-Instruct as the judge model and run it using bfloat16 layer-parallel inference over two H100 GPUs using HuggingFace transformers. We use prompts adapted from Li et al. (2023) and Villegas et al. (2025) for evaluating the general correctness of generated definitions as well as the similarity with the original model.
@@
B.7 Generation of Relevant Contexts
When generating contexts that contain a new token, we prompt the model with the beginning-of-sequence (BOS) token followed by the textual representation of the new token. Using the new token <palatable> as an example, our prompt is then: “<s> palatable”. Note that we add a prefix space after the BOS token <s>. We then sample new sequences of length , where and are the same as in the settings we use for context retrieval from an existing corpus (in this work we use and ). We sample using the standard sampling settings of each model from the transformers (Wolf et al., 2020) library.
B.8 Isolated Per-Token Optimization vs. Joint Optimization
When implementing Token Distillation, we have the choice of performing the optimization process for each new input embedding separately or simply allowing updates to all new input embeddings at the same time.
@@ B.9 Hyperparameters
We use AdamW (Kingma & Ba, 2017; Loshchilov & Hutter, 2019) optimization for all trainable parameters (the new token embeddings) with a batch size of 16, which maximizes throughput on our used Nvidia H100 80GB GPUs. We do not incorporate any weight decay and maintain a constant learning rate schedule with a linear warmup for the first half of the training steps. For fair comparison, we sweep for the best learning rate for each method that requires a learning rate. We use the same data for all training-based methods, including baselines and variations of our method.
B.10 Models
Appendix C Compute Budget & Runtime
We report the average training time for Token Distillation as well as NTP for initializing the new biomedical domain tokens per model in Table 10. We also report the number of new tokens per model. These differ slightly, as we build an initial fixed candidate list, which is then filtered against the existing vocabularies of each model.
Additional computational budget.
Since Token Distillation includes an additional forward pass through the model for generating “teacher” hidden states, we do see this overhead reflected in the run times. A forward pass can be approximated as taking one third of the computational cost of a complete forward-backward pass – which roughly matches our measured run times. Note however that – as demonstrated in Section 5.2 – Token Distillation can be significantly sped up while potentially even improving results by choosing an earlier target layer than the last one, yielding even faster run times than NTP.
However, our main version of Token Distillation is slightly more expensive than NTP – even though both have the same data budget. Thus, we also run the baseline NTP with an additional epoch to compensate for the additional teacher forward pass that Token Distillation uses. Note that this in fact significantly overcompensates for the additional budget Token Distillation uses. We report the results in Table 11. NTP is not able to take any major advantage of an additional epoch, yielding similar results to using a single epoch. Investigating the loss curves, we find that the training converges towards the beginning of the second epoch.555Note that we do not anneal the learning rate to zero, so this is not an artifact of the learning rate schedule.
We believe that this is at least partly due to the fundamental limitation of next-token prediction for our task of learning a single new embedding for an already pretrained model. Instead of matching model behavior between seeing the single new and the multiple original subtokens, NTP simply adjusts weights to maximize the likelihood of the given sequences. Consider “<new_token> is a football player”, where we want <new_token> to represent <Messi>: NTP will only be able to learn a generic embedding from this sequence that will capture semantics from many different (American football and soccer) players. Our proposed distillation objective Token Distillation however is able to learn a more specific representation, as it has access to hidden states from the sequence “Messi is a football player”, which capture more specific details necessary for a good representation.
To demonstrate, we also run Token Distillation for an additional epoch. Token Distillation is in fact able to take advantage of a second pass over the data, in some case with major improvements in the benchmark results. This illustrates the richer training signal that our distillation-based objective provides. Remarkably, for OLMo2-7B-i, Token Distillation now is actually able to match the original model with the original tokenization while instead using the new tokenization.
Init Method Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i Llama3-8B Llama3-8B-i Mistral-7B OLMo2-7B-i NTP 08:44 08:49 05:52 05:18 08:43 08:46 06:01 07:40 Token Distillation 11:31 11:33 07:38 08:07 12:00 11:29 09:09 10:22 # new tokens 2589 2589 2589 2589 2589 2589 2637 2591
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Target 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 NTP (1 epoch) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 Token Distillation (1 epoch) 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 NTP (2 epoch) 61.0 66.1 65.4 65.9 68.7 56.7 61.5 58.9 63.0 Token Distillation (2 epoch) 63.2 68.5 68.9 67.7 71.8 58.4 63.6 61.3 65.4
Token Distillation versus hyper-networks.
Comparing the computational budget with ZeTT (Minixhofer et al., 2024), ZeTT sees 3.2 billion tokens during hyper-network pretraining, while Token Distillation sees only 3.2 million tokens (for initializing 2,600 new tokens), less than 1% of ZeTT’s budget – still Token Distillation outperforms ZeTT. In return, ZeTT is faster at inference time (in our experiments on a single H100 80GB GPU, ZeTT took less than a minute). In future work, we believe that investigating a Token Distillation-style objective for hyper-network pretraining would be beneficial, as our experiments show that this significantly outperforms next-token prediction for learning new embeddings (ZeTT uses next-token prediction).
A note on downstream computational speedups.
Over-tokenization is not only associated with reduced downstream task performance (Rust et al., 2021; Ali et al., 2024) but also with increased computational cost due to longer sequence lengths (Ahia et al., 2023; Yamaguchi et al., 2024a). This computational cost is reduced when reducing over-tokenization by adding new (commonly occurring) tokens to the vocabulary. Importantly, for input embeddings, the reduction in computational cost is completely independent of the actual quality of the embeddings for the newly added tokens (for a given fixed sequence of text). The same computational speedups are achieved by using Random initialization and much better methods such as ZeTT (Minixhofer et al., 2024) or our proposed Token Distillation. Of course, the actual quality of any predictions made conditioned on a sequence containing such new tokens does heavily depend on the embedding quality. Therefore, in this work, we focus on direct evaluation of the quality of produced embeddings.